# Panoramic Distortion-Aware Tokenization for Person Detection and Localization Using Transformers in Overhead Fisheye Images

### v2

@@ Abstract
Person detection methods are used widely in applications including visual surveillance, pedestrian detection, and robotics. However, accurate detection of persons from overhead fisheye images remains an open challenge because of factors including person rotation and small-sized persons. To address the person rotation problems, we convert the fisheye images into panoramic images. For smaller people, we focused on the geometry of the panoramas. Conventional detection methods tend to focus on larger people because these larger people yield large significant areas for feature maps. In equirectangular panoramic images, we find that a person’s height decreases linearly near the top of the images. Using this finding, we leverage the significance values and aggregate tokens that are sorted based on these values to balance the significant areas. In this leveraging process, we introduce panoramic distortion-aware tokenization. This tokenization procedure divides a panoramic image using self-similarity figures that enable determination of optimal divisions without gaps, and we leverage the maximum significance values in each tile of the token groups to preserve the significant areas of smaller people. To achieve higher detection accuracy, we propose a person detection and localization method that combines panoramic-image remapping and the tokenization procedure. Extensive experiments demonstrated that our method outperforms conventional methods when applied to large-scale datasets.
@@ 1 Introduction
Figure 1: Concept illustrations of our work. Our method converts an overhead fisheye image into an equirectangular panoramic image by remapping. To leverage the significant areas of smaller people, a distortion-aware tokenization process based on self-similarity figures using a unit figure, which is colored dark red and consists of a square and two rectangles. Our transformer-based detector uses the resulting tokens to detect person-bounding boxes within the panoramic image. We then obtain the bounding boxes in the fisheye images from the panoramic image via projection. The input image is taken from [64].
@@
Object detection methods are used in a wide range of applications in both the commercial and industrial fields, including visual surveillance [11, 23, 62, 67], pedestrian detection [7, 19, 69], and robotics [21, 65, 68]. In particular, the interest in fisheye image-based person detection has been growing because of its varied applications, which include monitoring of public transportation and infrastructure, worker movement optimization in factories, and player movement analysis for use in sports strategy. Overhead fisheye cameras are suitable for use in these applications because they are omnidirectional and have a larger field of view than conventional narrow-view cameras; i.e., one overhead fisheye camera can substitute for several narrow-view cameras. However, the overhead fisheye images do not match the inputs of the detectors [13, 59, 73], which require perspective images. This mismatch is derived from person rotation and from the presence of small-sized people in the images. The person rotation factor leads to individual persons having various shapes and appearances in the fisheye images. Within the near image circle of circumferential fisheye images, smaller people are captured at a long distance from the fisheye cameras. Although person detection methods based on use of narrow-view cameras are well-established, deriving accurate techniques for person detection and localization from overhead fisheye images has remained an open challenge.
@@
Person detection methods have been proposed that used convolutional neural networks (CNNs) [20] and transformers [53] in overhead fisheye images. Early CNN-based methods [22, 41] used detectors that were trained on perspective images. To improve the performance of the CNN-based methods [10, 34, 47] and a transformer-based method [64], their training was also conducted using overhead fisheye images. However, it is difficult for these conventional methods to detect people accurately in overhead fisheye images because of the rotation and small person size problems mentioned above.
@@
On the basis of the observations above, to realize accurate person detection, we propose a transformer-based method that detects rotated bounding boxes from an overhead fisheye image, as illustrated in Figure 1. An input fisheye image is converted into a panoramic image using a remapping method. For the panoramic image, we introduce a tokenization method called panoramic distortion-aware tokenization (PDAT) to balance the significant areas of feature maps. This tokenization uses self-similarity figures to perform nonuniform divisions by matching the token tile size with a person’s height; i.e., the height depends on the vertical image coordinates of the panoramic image. Finally, we obtain bounding boxes in the fisheye images via projection.
@@
To investigate the effectiveness of these proposed methods, we conducted extensive experiments on three large-scale datasets[10, 48, 64]. This evaluation demonstrated that our proposed method outperforms both the conventional CNN-based [10, 22, 34, 41, 47] and transformer-based [64] methods. The major contributions of our study are summarized as follows:
@@
  • We propose a transformer-based method for person detection and localization in overhead fisheye images to achieve higher accuracy than the conventional methods without the need to balance significant areas of the feature maps.
@@
  • We introduce the PDAT method with optimal divisions based on self-similarity figures to address a person’s height using the vertical panoramic-image coordinates.
@@
  • We demonstrate the effectiveness of our proposed method by applying it to the challenging LOAF dataset. The average precision for distant person positions when using our method on the LOAF test set is substantially greater than that in the existing state-of-the-art method by 32.0.
@@ 2 Related work
Core idea to address
Method Detector 1 Training on 2 Bounding box Person Small
OH fisheye rotation person
CNN-based
Seidel et al. [41] VISAPP’19 YOLOv2 [35] Orthogonal rectangle
Li et al. [22] AVSS’19 YOLOv3 [36] Rotated rectangle
Tamura et al. [47] WACV’19 YOLOv2 [35] Rotated rectangle
RAPiD [10] CVPRW’20 FPN [25] Rotated rectangle
OARPD [34] MTA’24 YOLOv8m [52] Rotated rectangle
Transformer-based
Yang et al. [64] ICCV’23    DAB-DETR [27] Rotated rectangle
Ours DAB-DETR [27] Rotated trapezoid       ✓       ✓
2 Networks were trained on overhead fisheye images.
@@
Overhead-fisheye-image person detection. Person detection techniques in overhead fisheye images have been developed using various datasets [5, 10, 22, 48, 64]. Existing person detection methods for overhead fisheye images can be classified into two categories: CNN-based and transformer-based methods, as listed in Table 1. Seidel et al. [41] proposed a pioneering CNN-based detection method. The method divides a fisheye image into upright image patches that are undistorted using the given camera parameters. Although these patches are perspective images, the method requires camera calibration before detection. Li et al. [22] proposed a CNN-based method to improve detection accuracy using focal windows that can extract upright images. Tamura et al. [47] proposed a method that uses YOLOv2 [35], which had been trained on perspective images from the COCO dataset [24] using rotation augmentation. To train a detector called RAPiD, Duan et al. [10] introduced the angle-aware loss function. Inspired by the RAPiD detector being trained on fisheye images, OARPD [34] was proposed with use of the center distance intersection over union approach to address rotated bounding boxes. Yang et al. [64] proposed a pioneering transformer-based method with rotation-equivariant training. For a query-based detector, this training strategy involves use of an additional query of a $90^{\circ}$-rotated image to tackle rotated persons. Both the CNN-based and transformer-based methods described above focused primarily on person rotation. Although the CNN-based and transformer-based methods can detect people, images containing people of various sizes cause performance degradation.
Panoramic-image conversion. When the camera parameters are given, the panoramic images, fisheye images, and spherical representations are all convertible. This conversion method can be applied in various tasks using fisheye and panoramic images. To alleviate fisheye distortion, Plaut et al. [33] converted fisheye images into equirectangular images to perform 3D object detection. In deep single image camera calibration, equirectangular images were converted into fisheye images for dataset generation [29, 55]. Wang et al. [57] used spherical representations taken from equirectangular images to perform 360 room layout estimation. As described in the discussion of overhead-fisheye-image person detection above, conventional methods do not convert the fisheye images into panoramic images because person detection in panoramic images remains an open challenge.
@@
Small object detection. Object detection accuracy tends to decrease for smaller objects. Small objects are often captured at long distances, e.g., in a dataset acquired using drones [72]. Small object detection methods are generally classified into three categories: division-based methods, augmentation-based methods, and enlargement-based methods. The division-based methods involve multi-scale training performed using image pyramids [44], image cropping [45], and feature pyramids [14]. The augmentation-based methods [13, 66] were proposed to realize data invariance. The enlargement-based methods enlarge the input images [26, 51, 58] and conduct super-resolution processing [17, 43]. The methods noted above do not use the prior knowledge that smaller people are only located near the image circles in overhead fisheye images. Therefore, these conventional methods do not work effectively for person detection in overhead fisheye images.
Visual localization. Various localization methods have been established to determine either camera poses or person locations from an image. Visual localization methods can be classified into three categories: query-based methods, regression-based methods, and geometry-based methods. The query-based methods estimate camera poses using databases composed of geo-tagged images [1, 3, 16, 40, 49, 63] and 3D scene structure models [12, 39]. Before query-based methods can be used, the databases and models must first be built in a time-consuming process. The regression-based methods [18, 56] can estimate camera poses without the need for databases and models; however, their inferences can only be applied to the same scenes that were used for training. The geometry-based methods estimate person positions based on the bounding boxes [64, 70] and head points [6] in overhead fisheye images. The geometry-based method is described in detail below.
Yang et al.’s localization method [64] estimates person locations based on the camera geometry after camera calibration. The camera models express the required mapping from world coordinates $\tilde{\mathbf{p}}$ to image coordinates $\tilde{\mathbf{u}}$ in homogeneous coordinates. This mapping is represented as
$\tilde{\mathbf{u}}=\left[\begin{array}[]{ccc}\gamma/d_{u}&0&c_{u}\\ 0&\gamma/d_{v}&c_{v}\\ 0&0&1\end{array}\right]\left[~{}\mathbf{R}~{}|~{}\mathbf{t}~{}\right]\tilde{\mathbf{p}},$ (4)
where $\gamma$ is the distortion, $\left(d_{u},d_{v}\right)$ is the image sensor pitch, $\left(c_{u},c_{v}\right)$ is a principal point, $\mathbf{R}$ is a rotation matrix, and $\mathbf{t}$ is a translation vector. The subscripts $u$ and $v$ denote the horizontal and vertical directions, respectively. Given the camera parameters used in Equation (4), we can determine the 3D vector of the incident ray from $\tilde{\mathbf{u}}$. In the following explanation, we define the incident angle $\theta$ as being formed between the optical axis and the incident ray. The line that passes through a camera along the incident ray vector crosses the ground plane. Therefore, we can calculate the distance between the crossing point and the ground point beneath the camera for a person when $\tilde{\mathbf{u}}$ is the center of the lower edge of a detected person-bounding box. However, accurate person detection is essential because the estimated location is sensitive to person detection errors. We therefore focused on geometry-based localization; i.e., the process is adapted to surveillance using overhead fisheye cameras.
@@
Transformer tokenization. Dosovitskiy et al. [9] proposed a pioneering transformer-based method using a tokenization approach that divides an input image into square patches. Numerous tokenization methods have been proposed to date because these square patches lead to high computational costs. Window-based methods focus on the token groups using shifted windows [28], cross-shaped windows [8], dilated windows [50, 60], and scalable sliding windows [15]. In contrast to static window shapes, deformable attention methods [61] were proposed to improve flexibility. To address the problem of multiple tokens being generated in unimportant regions, e.g., in the background, BiFormer [71] prunes the regions using a coarse-to-fine strategy and SG-Former [37] aggregates the regions. Although the methods described above are all well-developed, these methods were designed for use with perspective images. For fisheye image applications, DarSwin [2], which uses radial patches formed by polar partitioning, was proposed. However, this partitioning approach is not suitable for person detection in overhead fisheye images because a small number of the relevant tokens are generated in near image circles.
@@ 3 Proposed method
In this section, we first analyze the bounding boxes in an overhead fisheye dataset. Second, we introduce PDAT and depict how it groups the tokens. Third, we describe our transformer-based method for person detection and localization using PDAT. Finally, we describe our training and inference phases.
3.1 Analysis of an overhead fisheye dataset
The geometric relationship between equirectangular images and overhead fisheye cameras leads to the fact that the heights and widths of the bounding boxes decrease linearly when the incident angles $\theta$, defined in Section 2 (visual localization), are large enough. We verified the relationship using the LOAF training set. This distribution indicates that the incident angles that are greater than approximately 70 satisfy the required relationship. In addition, the range of incident angles from 80 to 90 dominated 75% of the number of bounding boxes. Therefore, we design our tokenization method based on the relationship above, which we explain in further detail in the following.

### v3

# Panoramic Distortion-Aware Tokenization for Person Detection and Localization in Overhead Fisheye Images

@@ Abstract
Person detection in overhead fisheye images is challenging due to person rotation and small persons. Prior work has mainly addressed person rotation, leaving the small-person problem underexplored. We remap fisheye images to equirectangular panoramas to handle rotation and exploit panoramic geometry to handle small persons more effectively. Conventional detection methods tend to favor larger persons because they dominate the attention maps, causing smaller persons to be missed. In hemispherical equirectangular panoramas, we find that apparent person height decreases approximately linearly with the vertical angle near the top of the image. Using this finding, we introduce panoramic distortion-aware tokenization to enhance the detection of small persons. This tokenization procedure divides panoramic features using self-similar figures that enable the determination of optimal divisions without gaps, and we leverage the maximum significance values in each tile of the token groups to preserve the significance areas of smaller persons. We propose a transformer-based person detection and localization method that combines panoramic-image remapping and the tokenization procedure. Extensive experiments demonstrated that our method outperforms conventional methods on large-scale datasets.
@@ 1 Introduction
Figure 1: Our method converts an overhead fisheye image into an equirectangular image by remapping. To leverage the significance areas of smaller persons, a distortion-aware tokenization process based on a self-similar figure using a unit figure, which is colored dark red and consists of a square and two rectangles. Our transformer-based detector uses the resulting tokens to detect bounding boxes within the panoramic image, and then the boxes project to the fisheye image. The input image is taken from [Yang_ICCV_2023].
Method Detector 1 Training on overhead fisheye images Bounding box Core idea to address
 Person rotation    Small person
CNN-based
Seidel et al[Seidel_VISAPP_2019] VISAPP’19 YOLOv2 [Redmon_CVPR_2017]     Orthogonal rectangle
Li et al[Li_AVSS_2019] AVSS’19 YOLOv3 [Redmon_arXiv_2018] Rotated rectangle
Tamura et al[Tamura_WACV_2019] WACV’19 YOLOv2 [Redmon_CVPR_2017] Rotated rectangle
RAPiD [Duan_CVPRW_2020] CVPRW’20 FPN [Lin_CVPR_2017] Rotated rectangle
OARPD [Qiao_MTA_2024] MTA’24 YOLOv8m [Varghese_ADICS_2024] Rotated rectangle
Transformer-based
Yang et al[Yang_ICCV_2023] ICCV’23    DAB-DETR [Liu_ICLR_2022] Rotated rectangle
Ours DAB-DETR [Liu_ICLR_2022] Rotated trapezoid       ✓       ✓
@@
Object detection methods are used in a wide range of applications in both the commercial and industrial fields, including visual surveillance [Fu_CVPR_2021, Lin_TIP_2012, Xie_ICPR_2010, Yuan_CVPR_2024], pedestrian detection [Dollar_CVPR_2009, Khan_CVPR_2023, Zhang_PAMI_2018], and robotics [Lee_ICRA_2018, Yoshimi_IROS_2006, Zhang_IROS_2024]. In particular, the interest in fisheye image-based person detection has been growing because of its varied applications, which include monitoring of public transportation and infrastructure, worker movement optimization in factories, and player movement analysis for use in sports strategy. Overhead fisheye cameras are suitable for use in these applications because they are omnidirectional and have a larger field of view than conventional narrow-view cameras; i.e., one overhead fisheye camera can substitute for several narrow-view cameras. However, the overhead fisheye images do not match the inputs of the detectors [Ghiasi_CVPR_2021, Huang_2025_CVPR, Wang_CVPR_2023, Zhang_2025_CVPR, Zong_ICCV_2023], which require perspective images. This mismatch is derived from person rotation and from the presence of small persons in the images. The person rotation factor leads to individual persons having various shapes and appearances in the fisheye images. Within the near image circle of circumferential fisheye images, smaller people are captured at a long distance from the fisheye cameras. Although person detection methods based on use of narrow-view cameras are well-established, deriving accurate techniques for person detection and localization from overhead fisheye images has remained an open challenge.
@@
Person detection methods have been proposed that used convolutional neural networks (CNNs) [Krizhevsky_NIPS_2012] and transformers [Vaswani_NIPS_2017] in overhead fisheye images. Early CNN-based methods [Li_AVSS_2019, Seidel_VISAPP_2019] used detectors that were trained on perspective images. To improve the performance of the CNN-based methods [Duan_CVPRW_2020, Qiao_MTA_2024, Tamura_WACV_2019] and a transformer-based method [Yang_ICCV_2023], their training was also conducted using overhead fisheye images. However, it is difficult for these methods to detect persons accurately in overhead fisheye images because of the rotation and small-person problems.
@@
On the basis of the observations above, to realize accurate person detection, we propose a transformer-based method that detects rotated bounding boxes from an overhead fisheye image, as illustrated in Figure 1. An input fisheye image is converted into a panoramic image using a remapping method. For the panoramic image, we introduce a tokenization method called panoramic distortion-aware tokenization (PDAT) to balance the significance areas of feature maps regardless of person sizes. This tokenization uses self-similar figures to perform nonuniform divisions while maintaining consistent person-to-tile size ratios, based on our finding that person height depends on the vertical image coordinates of the panoramic image. Finally, we obtain bounding boxes in the fisheye images via projection.
@@
To investigate the effectiveness of these proposed methods, we conducted extensive experiments on three large-scale datasets [Duan_CVPRW_2020, Tezcan_WACV_2022, Yang_ICCV_2023]. This evaluation demonstrated that our proposed method outperforms both the conventional CNN-based [Duan_CVPRW_2020, Li_AVSS_2019, Qiao_MTA_2024, Seidel_VISAPP_2019, Tamura_WACV_2019] and transformer-based [Yang_ICCV_2023] methods. The major contributions of our study are summarized as follows:
@@
  • We propose a transformer-based method for person detection and localization in overhead fisheye images that achieves higher accuracy than conventional methods by balancing significance areas of the feature maps.
@@
  • We introduce the PDAT method with optimal divisions based on self-similar figures to address a person’s height using the vertical panoramic-image coordinates.
@@
  • We demonstrate the effectiveness of our proposed method by applying it to the challenging LOAF dataset. The average precision for distant person positions when using our method on the LOAF test set is substantially greater than that of the existing state-of-the-art method by 32.0.
@@ 2 Related work
Overhead-fisheye-image person detection. Person detection techniques in overhead fisheye images have been developed using various datasets [Blanco_SPIC_2021, Duan_CVPRW_2020, Li_AVSS_2019, Tezcan_WACV_2022, Yang_ICCV_2023]. Existing person detection methods for overhead fisheye images can be classified into two categories: CNN-based and transformer-based methods, as listed in Table 1. Seidel et al[Seidel_VISAPP_2019] proposed a pioneering CNN-based detection method. The method divides a fisheye image into upright image patches that are undistorted using the given camera parameters. Although these patches are perspective images, the method requires camera calibration before detection. Li et al[Li_AVSS_2019] proposed a CNN-based method to improve detection accuracy using focal windows that can extract upright images. Tamura et al[Tamura_WACV_2019] proposed a method that uses YOLOv2 [Redmon_CVPR_2017], which had been trained on perspective images from the COCO dataset [Lin_ECCV_2014] using rotation augmentation. To train a detector called RAPiD, Duan et al[Duan_CVPRW_2020] introduced the angle-aware loss function. Inspired by RAPiD being trained on fisheye images, OARPD [Qiao_MTA_2024] was proposed with use of the center distance intersection over union approach to address rotated bounding boxes. Yang et al[Yang_ICCV_2023] proposed a pioneering transformer-based method with rotation-equivariant training. For a query-based detector, this training strategy involves use of an additional query of a $90^{\circ}$-rotated image to tackle rotated persons. Both the CNN-based and transformer-based methods described above focused primarily on person rotation. Although the CNN-based and transformer-based methods can detect persons, images containing persons of various sizes cause performance degradation.
@@
Small object detection. Object detection accuracy tends to decrease for smaller objects. Small objects are often captured at long distances, e.g., in a dataset acquired using drones [Zhu_arXiv_2018]. Small object detection methods are generally classified into four categories: division-based methods, augmentation-based methods, enlargement-based methods, and spectrum-based methods. The division-based methods involve multi-scale training performed using image pyramids [Singh_CVPR_2018], image cropping [Singh_NIPS_2018], and feature pyramids [Gong_WACV_2021]. The augmentation-based methods [Ghiasi_CVPR_2021, Yu_WACV_2020] were proposed to realize data invariance. The enlargement-based methods enlarge the input images [Liu_TIP_2025, Unel_CVPRW_2019, Wang_PAMI_2021] and conduct super-resolution processing [Ji_ICPR_2021, Shermeyer_CVPRW_2019]. The spectrum-based methods enhance the discrimination of objects from a frequency perspective [Sun_2025_CVPR]. The methods noted above do not use the prior knowledge that smaller persons are only located near the image circles in overhead fisheye images. Therefore, these conventional methods do not work effectively for person detection in overhead fisheye images.
@@
Transformer tokenization. Dosovitskiy et al[Dosovitskiy_ICLR_2021] proposed a pioneering transformer-based method using a tokenization approach that divides an input image into square patches. Numerous tokenization methods have been proposed to date because these square patches lead to high computational costs. Window-based methods focus on the token groups using shifted windows [Liu_ICCV_2021], cross-shaped windows [Dong_CVPR_2022], dilated windows [Tu_ECCV_2022, Wang_PAMI_2024], and scalable sliding windows [Hassani_CVPR_2023]. In contrast to static window shapes, deformable attention methods [Xia_CVPR_2022] were proposed to improve flexibility. To address the problem of multiple tokens being generated in unimportant areas, e.g., in the background, DynamicViT [Rao_NIPS_2021] and BiFormer [Zhu_CVPR_2023] prune the areas, and SG-Former [Ren_ICCV_2023] aggregates the areas. Although the methods described above are all well-developed, these methods were designed for use with perspective images. For fisheye image applications, DarSwin [Athwale_ICCV_2023], which uses radial patches formed by polar partitioning, was proposed. However, this partitioning approach is not suitable for overhead fisheye images because a small number of the relevant tokens are generated in near image circles.
Visual localization. Various localization methods have been established to determine either camera poses or person locations from an image. These methods can be classified into three categories: query-based methods, regression-based methods, and geometry-based methods. The query-based methods estimate camera poses using databases composed of geo-tagged images [Aiger_ICCV_2023, Balntas_ECCV_2018, Hyeon_ICCV_2021, Sarlin_CVPR_2021, Torii_PAMI_2021, YangL_ICCV_2019] and 3D scene structure models [Germain_CVPR_2021, Sarlin_CVPR_2019]. Before query-based methods can be used, the databases and models must first be built in a time-consuming process. The regression-based methods [Kendall_ICCV_2015, Walch_ICCV_2017] can estimate camera poses without the need for databases and models; however, their inferences can only be applied to the same scenes that were used for training. The geometry-based methods estimate horizontal distances of person positions based on the bounding boxes [Yang_ICCV_2023, Zhu_JVCIR_2019] and head points [DelBlanco_SPIC_2021] in overhead fisheye images. These geometry-based methods estimate person locations from image coordinates using camera parameters.
[…]
Figure 3: Panoramic distortion-aware tokenization. (a) A panoramic feature map is nonuniformly tiled with cyan grids along the incident-angle axis. The tile sizes are adjusted to match the person’s height in each region, maintaining similar person-to-tile size ratios across different regions. Enlarged views show person-bounding boxes in regions 3, 4, and 5. (b) A self-similar subdivision recursively partitions the unit figure and is truncated after three steps, yielding five regions ($K=5$) with the size ratios of the unit figure: 1, 1/2, 1/4, and 1/8.
@@ 3 Proposed method
First, we analyze the bounding boxes in an overhead fisheye image dataset. Second, we present the design of our method. Third, we introduce PDAT. Fourth, we describe our method using PDAT for person detection and localization. Finally, we describe our training and inference phases.
3.1 Geometric analysis of overhead fisheye images
We analyzed the LOAF training set [Yang_ICCV_2023] to understand the unique characteristics of overhead fisheye images, which differ from perspective images. We remapped overhead fisheye images to equirectangular images, as shown in Figure 2(a). A downward-looking fisheye camera observes the lower hemisphere, with incident angles $\theta$ ranging from 0 (nadir) to 90 (horizon). Figure 2(b) shows the distribution of bounding box heights and widths along different incident angles $\theta$. Our analysis revealed two key findings:
  • We found a geometric relationship between equirectangular images and overhead fisheye cameras: the heights and widths of person-bounding boxes decrease approximately linearly when the incident angles $\theta$ are large enough ($70^{\circ}$).
  • We observed that 75% of all person-bounding boxes are concentrated in the range of incident angles from 80 to 90, indicating a strong bias in person distribution.
This pattern in the first finding arises from the geometry of overhead fisheye cameras and is fundamentally different from that observed with perspective cameras, where object sizes primarily depend on distance. While the heights at smaller incident angles exhibit an upward-convex trend due to oblique viewing angles, we focus on the linear behavior at large incident angles, where the majority of bounding boxes are located, as indicated in the second finding.
Note that the bias in the second finding has two main causes. First, peripheral portions at large incident angles subtend larger ground areas. Second, due to motion parallax [Gibson_JEP_1959, Hartley_CUP_2004], people farther from the camera exhibit smaller apparent image motion and remain in view longer, while nearer people traverse the field of view more quickly.
3.2 Design of our proposed method
We designed our proposed method based on the two key findings, as described in Section 3.1, because conventional detection methods, which assume uniform object distribution, may not be optimal for overhead fisheye images. The first finding suggests that our method uses the prior knowledge of the linear change in predictable bounding box sizes. The second finding indicates that focusing on bounding boxes with large incident angles is effective in terms of the number of bounding boxes. Conventional detectors allocate computation approximately uniformly across the image and do not explicitly exploit $\theta$-dependent scale or distribution biases. Therefore, we designed our tokenization method to specifically account for these geometric relationships, focusing on the portions with large incident angles.